OpenStudy (anonymous):

A metal washer 1 cm in diameter is pierced by a 1/2 cm hole. What is the area of one face of the washer?

OpenStudy (jack1):

ok, where you wwant to start is work out the area of the washer... before it had the hole cut in it so it's just the equation for area of a circle, are u right with that?

OpenStudy (anonymous):

Nope.

OpenStudy (jack1):

ok... area of a circle = pi x (radius)^2 your radius is half the length of the diameter (which was 1 cm) so whats the area equal?

OpenStudy (anonymous):

Pi/4

OpenStudy (jack1):

cool now subtract the area of the hole from the area of the washer Area of hole = pi x (radius squared) radius of hole = 1/4 cm (given diameter of hole =1/2 cm

OpenStudy (jack1):

so what did you get for area of the hole?

OpenStudy (anonymous):

pi/16

OpenStudy (nurali):

area = pi*r^2 = pi*(1/2)^2 = pi*1/4 next, calculate the center hole of diameter 1/2 area :- area = pi*r^2 = pi*(1/4)^2 = pi*1/16 area of washer face = big circle area - hole area = pi*1/4 - pi*1/16 = pi*4/16 - pi*1/16 = pi(4-1)/16 = 3pi/16

OpenStudy (jack1):

yep, so pi/4 - pi/16 = area that's left = 3pi/16 cm^2 or 0.1875 pi cm^2 or 0.589 cm^2

OpenStudy (anonymous):

Aah, okay. Thank You :D

OpenStudy (anonymous):

@Jack1 Two spheres of copper, of radii 1cm and 2 cm, respectively, are melted inside a cylinder of radius 1 cm. Find the altitude of the cylinder.

OpenStudy (jack1):

...altitude... height above the ground...? what do you mean? ...how full is the cylinder in terms of height?

OpenStudy (nurali):

vol of sphere 1 = (4/3)*pi*1^3 vol of sphere 2 = (4/3)*pi*2^3 total volume of melted material = (4/3)*pi*9 ( sum of the above) vol of the moulded cone = (1/3)*pi*r^2*h, where r = radius and h is height. Given r=1 So we have (4/3)*pi*9 = (1/3)*pi*h; cancelling pi on both sides, we have h = 36 cm

OpenStudy (anonymous):

I dunno. It's just my assignment.

OpenStudy (jack1):

@Nurali you type quick ;D

OpenStudy (anonymous):

But the choices are: a. 2 b. 3 c. 4 d. 12